OpenStudy (anonymous):

solve ??? 8(2 1/2x-3/4) - 1/4(1-x)=1/2

OpenStudy (anonymous):

What have you tried so far?

OpenStudy (anonymous):

I tried to distribute.

OpenStudy (anonymous):

What did you get when you distributed?

jhonyy9 (jhonyy9):

8(5x/2 -3/4) - (1-x)/4=1/2 8((10x-3)/4) -(1-x)/4 = 2/4 80x-24-1+x=2 81x=27 x=27/81 x=1/3